We have recovered the LN3 configuration. The recovery went smoothly, reached LN3 at the second attempt after one unlock in locking LN2.
We tested first the configuration in which the beam is centered in the West end mirror.
In figure 1, the comparison between the "after-cleaning" in purple and the "pre-cleaning" in blue sensitivities is shown (beam centered).
- the 3Hz is not present and the low frequency sensitivity is back to the best level
- the frequency noise at high frequency is much higher than before.
We left the ITF in LN3 for ~15 mins and we decided to unlock to test the decentered position of the beam on the WE (12mm offset).
We unlocked at locking in LN3, see next entry.
We leave in auto-relock with the decentering offset on.